Skip to main content

IT Status Alerts

All Closed Alerts

Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245

4882 results found, displaying results 101 - 121.

Performance Issue: Service restored: Off-campus phone calls to Cornell phones

Start:
2017-11-15 15:22:00
Status:
Closed
Brief Description:
*Service has been restored.* Due to a widespread service issue involving Verizon and AT&T, phone calls from off-campus numbers to Cornell (Ithaca campus) phone numbers were not able to be completed.
Current Status:
Phone service on the Ithaca campus has been fully restored. If you experience any issues with your Cornell phone service, please contact the IT Service Desk.
more details

Scheduled Service Change: Brief network connectivity outage, Goldwin Smith node, 11/17/17

Start:
2017-11-17 05:00:00
Status:
Closed
Brief Description:
Brief outage to network connectivity in Goldwin Smith node on Friday, November 17th between 5:00am and 7:00am.
more details

Scheduled Service Change: Brief network connectivity outage, Rhodes Node, 11/17/17

Start:
2017-11-17 05:00:00
Status:
Closed
Brief Description:
Brief connectivity outage due to work to complete cutting over the Rhodes node to the new Juniper router, Friday, November 17, 2017, between 5:00am and 7:00am.
more details

Unplanned Outage: IT services affected by issue in CIT Server Farm

Start:
2017-11-13 07:17:00
Status:
Closed
Brief Description:
A couple of IT services are still affected by the issue with the CIT Server Farm. CIT and the teams that manage those services are continuing to work on restoring them.
Current Status:
All services affected by this outage have been restored.
more details

Performance Issue: Cornell e-lists (Lyris) down

Start:
2017-11-13 08:00:00
Status:
Closed
Brief Description:
The main Cornell e-list (Lyris) service is down due to the issue with the CIT Server Farm. Messages sent to e-lists are being queued and will be sent once the service is available.
Current Status:
Lyris service has been restored. New messages are being delivered. Messages that were queued on Nov. 13 and Nov. 14 are being held back for now.
more details

Performance Issue: Office 365 email and Outlook on the Web (outlook.cornell.edu)

Start:
2017-11-14 08:30:00
Status:
Closed
Brief Description:
The ability to connect to Office 365 email and Outlook on the web (outlook.cornell.edu) is being interrupted while the server its authentication is on restarts.
Current Status:
Individuals should be able to connect reliably to Office 365 at this time. Engineers are continuing to monitor the situation. Anyone who has problems connecting should contact the IT Service Desk.
more details

Unplanned Outage: Student Essentials Unavailable

Start:
2017-11-13 08:00:00
Status:
Closed
Brief Description:
Student Essentials Site Unavailable
Current Status:
Student Essentials was affected by the CIT Server Farm issue. It is now back online.
more details

Scheduled Service Change: OBIEE Production unavailable 11/12/17 during maintenance

Start:
2017-11-12 06:00:00
Status:
Closed
Brief Description:
Oracle Business Intelligence Enterprise Edition (OBIEE) production will be unavailable during maintenance on Sunday, 11/12/17, from 6:00am - 7:00am.
more details

Unplanned Outage: Cornell SharePoint Online now available

Start:
2017-10-30 12:11:00
Status:
Closed
Brief Description:
Cornell's SharePoint Online sites are now available.
Current Status:
Cornell's SharePoint Online sites are now available.
more details

Scheduled Service Change: Bi-weekly migration of PeopleSoft will cause an outage 11/09/2017

Start:
2017-11-09 05:00:00
Status:
Closed
Brief Description:
PeopleSoft will be unavailable on Thursday November 9, 2017 from approximately 5:00am to 7:00am.
more details

Security Alert: Large volume of spoofed email phishes

Start:
2017-10-19 13:30:00
Status:
Closed
Brief Description:
The IT Security Office is fielding an unusually large number of phishing emails with spoofed addresses that display the recipient in the from address.
Current Status:
The Security Office is still receiving reports of spoofed spam emails as of 11/09/17.
more details

Security Alert: Wireless WPA2 encryption vulnerability

Start:
2017-10-17 10:25:00
Status:
Closed
Brief Description:
A new critical vulnerability in secure Wi-Fi allows attackers to potentially view private information. Vendors are working to produce and distribute patches for affected devices. Patch as soon as possible and be vigilant about public wireless connections
more details

Performance Issue: Resolved Outlook on the web email messaging issue

Start:
2017-11-08 17:30:00
Status:
Closed
Brief Description:
There is currently an issue with Outlook Web App (OWA).
Current Status:
Microsoft has resolved the issue.
more details

Scheduled Service Change: Possible planned outage for Kronos after Daylight Saving Time on 11/5/17

Start:
2017-11-05 09:00:00
Status:
Closed
Brief Description:
On Sunday, November 5, 2017, it is possible Kronos will be unavailable to users beginning at 9am for up to 30 minutes. This planned outage may be needed to correct a potential problem with Kronos clocks after the Daylight Saving Time change.
more details

Scheduled Service Change: Move rhodesg3r1-5400 to new Rhodes Juniper router

Start:
2017-11-01 05:00:00
Status:
Closed
Brief Description:
Rhodes nodes switch impacting all campus VOIP.
more details

Scheduled Service Change: KFS, Cynergy, and Cynapps Update will cause outage 11/5/2017

Start:
2017-11-05 06:00:00
Status:
Closed
Brief Description:
KFS, eShop, Cynergy, Cynergy Applications and Travel will be down for new code deployment on Sunday, November 5th from 6am - Noon.
more details

Scheduled Service Change: KFS, e-Shop, Cynergy, and PCard Receipt will be unavailable 11/3/2017 6-7PM

Start:
2017-11-03 18:00:00
Status:
Closed
Brief Description:
Complete General Ledger Processing before tonight's scheduled 9:00 PM batch jobs. An ACL will be in place during this recovery effort to restrict access. Recovery efforts for failed GL Poster job in PROD from Thursday 11/02
more details

Performance Issue: Remedy Email Processing Delays

Start:
2017-11-06 08:53:00
Status:
Closed
Brief Description:
Email processing in Remedy is currently not working for all queues.
Current Status:
This issue has been resolved and email is now processing as expected in Remedy.
more details

Scheduled Service Change: CANCELLED - Entitlement changes and synchronization for 11/3/2017

Start:
2017-11-03 17:15:00
Status:
Closed
Brief Description:
CANCELLED - TO BE RESCHEDULED ... changes will be made to the Identity Management systems to complete the recent migration of employees from mail routing being processed internally to mail routing being processed by Office 365.
more details

Performance Issue: it.cornell.edu website issues

Start:
2017-11-01 14:24:00
Status:
Closed
Brief Description:
The it.cornell.edu website is intermittently going down. CIT is investigating the cause.
Current Status:
The instability with the it.cornell.edu website is no longer occurring.
more details
Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245